The difference of two integers is 9. Five times the smaller is 7 more than three times the larger. Find the numbers.
Answer by HyperBrain(694)

Let x=smaller no.
y=larger no.
y-x=9
[1]y=x+9
[2]5x=3y+3
Substituting [1] to [2],
5x=3(x+9)+3
5x=3x+27+3
5x=3x+30
2x=30
x=15
y-15=9
y=24
Therefore the two integers are 15 and 24.
Power up,
HyperBrain!
Power up,
HyperBrain!

Help us promote our advocacy:
http://uselessplastic.wordpress.com/
http://uselessplastic.multiply.com/